OpenStudy (anonymous):

what is the value of the expression 16/2^3 -1+3^3

OpenStudy (anonymous):

28

OpenStudy (anonymous):

16/(8) - 1 + (27) 2 - 1 + 27 = 28 PEMDAS - order of operations

OpenStudy (valpey):

If you mean: \[\frac{16}{2^{3}}-1+3^{3} = \frac{16}{8} - 1 + 27 = 2-1+27=28\]